Verify if the ISBN 978-0-760-73261-6 is valid?

1
Expert's answer
2022-03-09T14:56:29-0500

Solution


Given that ISBN 978-0-760-73261-6


The given ISBN is 13 digit ISBN number


We assign the position of each number from right to left


For

9 the position is 13

7 the position is 12

8 the position is 11

0 the position is 10

7 the position is 9

6 the position is 8

0 the position is 7

7 the position is 6

3 the position is 5

2 the position is 4

6 the position is 3

1 the position is 2

6 the position is 0


Now we the Odd number positions by 1 and Even number positions by 3.


"9 × (1) + 7 ×(3)+8 ×(1)+0 ×(3)+7 ×(1)+6 ×(3)+0 ×(1)+7 ×(3)+3 ×(1)+2 ×(3)+6 ×(1)+1 ×(3)+6 ×(1)\\
=108"



Now we divide this sum by "10" and see that the remainder is "8\u22600"


"\frac{108}{10}=10+\frac{8}{10}"


since the remainder is not zero, therefore, the given ISBN number is not a valid ISBN number.